# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ 6 Conclusion
In this paper, we propose PRISM, a robust skill-based meta-RL framework designed to address noisy offline demonstrations in long-horizon tasks. Through prioritized skill refinement and maximum return relabeling, PRISM effectively prioritizes task-relevant trajectories for skill learning and discovery. Experimental results highlight its robustness to noise and superior performance, demonstrating its potential for efficient meta-RL in real-world applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here
Appendix A Implementation Details on PRISM
This section provides a detailed implementation of the proposed PRISM framework. As outlined in Section 4, PRISM begins with an initial skill learning phase (pre-train) to train the low-level skill policy, skill encoder, and skill prior. It then progresses to the meta-train phase, where skill model refinement is performed through skill exploration using the exploration policy, and skill-based meta-RL is executed using the high-level policy and task encoder. Finally, in the meta-test phase, rapid adaptation to the target task is achieved via fine-tuning based on the trained high-level policy and task encoder. Section A.1 details the initial skill learning phase, Section A.2 elaborates on the meta-train phase, and Section A.3 explains the meta-test phase. All loss functions in PRISM are redefined in terms of the neural network parameters of its policies and models. Additionally, the overall structure for the meta-train and meta-test phases is provided in Algorithms 1 and 2.
A.1 Initial Skill Learning Phase
@@
$\mathcal{L}_{\text{spirl}}(\phi)$ $:=\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}_{\text{off}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl{[}\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr{]}$ (A.1)
@@
$=\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}_{\text{off}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Biggl{[}-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l,\phi}(a_{k}|s_{k},z)&+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\Bigl{(}q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\,\Big{|}\Big{|}\,\mathcal{N}(\mathbf{0},\mathbf{I})\Bigr{)}\\ &+\mathcal{D}_{\text{KL}}\Bigl{(}\lfloor q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\rfloor\,\Big{|}\Big{|}\,p_{\phi}(\cdot|s_{t})\Bigr{)}\Biggr{]},\end{aligned}$
@@
A.2 Meta-Train Phase
As described in Section 4, PRISM comprises three main processes: Skill Exploration, which explores diverse skills near the prioritized on-policy buffer $\mathcal{B}_{\text{on}}$; Prioritized Skill Refinement, which improves skills using $\mathcal{B}_{\text{on}}$; and Skill-based Meta-RL, which trains the high-level policy and task encoder to effectively utilize learned skills for solving tasks. Detailed explanations for each process are provided below.
Skill Exploration
As described in Section 4.2, the exploration policy $\pi_{\text{exp},\psi}$, parameterized by $\psi$, is designed to expand the skill distribution and discover task-relevant behaviors near trajectories stored in the prioritized on-policy buffer $\mathcal{B}_{\text{on}}^{i}$ for each training task $\mathcal{T}^{i}$. This buffer prioritizes trajectories that best solve the tasks. Additionally, the state-action value function $Q_{\text{exp},\psi}$, also parameterized by $\psi$, is defined to train the exploration policy using soft actor-critic (SAC). To enhance exploration, both extrinsic reward $r_{t}^{i}$ and intrinsic reward $r_{\text{int},t}^{i}$ are employed. The intrinsic reward, based on random network distillation (RND), is computed as the L2 loss between a randomly initialized target network $\hat{f}_{\bar{\eta}}^{i}$ and a prediction network $f_{\eta}^{i}$, parameterized by $\eta$ and $\bar{\eta}$, respectively, and is expressed as:
$r_{\text{int},t}^{i}:=\left\|f_{\eta}^{i}(s_{t+1})-\hat{f}_{\bar{\eta}}^{i}(s_{t+1})\right\|^{2}_{2},$ (A.2)
where $i$ is the task index, and $f_{\eta}$ is updated to minimize this loss. A dropout layer is applied to $f_{\eta}$ to prevent over-sensitivity to state $s$. The RL loss functions of SAC for training the exploration policy $\pi_{\text{exp},\psi}$ and the state-action value function $Q_{\text{exp},\psi}$ using the intrinsic reward $r_{\text{int},t}$ are defined as follows:
$\mathcal{L}_{\text{exp}}^{\text{critic}}(\psi):=\sum_{i}\mathbb{E}_{\begin{subarray}{c}(s_{t},a_{t},r_{t}^{i},s_{t+1})\sim\mathcal{B}^{i}_{\text{exp}}\cup\mathcal{B}^{i}_{\text{on}}\\ a_{t+1}\sim\pi_{\text{exp},\psi}(\cdot|s_{t},i)\end{subarray}}\biggl{[}\frac{1}{2}\biggl{(}Q_{\text{exp},\psi}(s_{t},a_{t},i)-\Bigl{(}\delta_{\text{ext}}r_{t}^{i}+&\delta_{\text{int}}r_{\text{int,t}}^{i}+\gamma_{\text{exp}}\bigl{(}Q_{\text{exp},\psi}(s_{t+1},a_{t+1},i)\\ &+\lambda^{\text{ent}}_{\text{exp}}\log\pi_{\text{exp},\psi}(a_{t+1}|s_{t+1},i)\bigr{)}\Bigr{)}\biggr{)}^{2}\biggr{]}\end{aligned}\\ &\begin{aligned} \mathcal{L}_{\text{exp}}^{\text{actor}}(\psi):=&\sum_{i}\mathbb{E}_{\begin{subarray}{c}s_{t}\sim\mathcal{B}^{i}_{\text{exp}}\cup\mathcal{B}^{i}_{\text{on}}\\ a_{t}\sim\pi_{\text{exp},\psi}(\cdot|s_{t},i)\end{subarray}}\biggl{[}\lambda^{\text{ent}}_{\text{exp}}\log\pi_{\text{exp},\psi}(a_{t}|s_{t},i)-Q_{\text{exp},\psi}(s_{t},a_{t},i)\biggr{]}-\lambda^{\text{kld}}_{\text{exp}}\sum_{i}\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{B}^{i}_{\text{on}}}\biggl{[}\log\pi_{\text{exp},\psi}(a_{t}|s_{t},i)\biggr{]}.\end{aligned}\end{aligned}$ (A.3)
Here, $\delta_{\text{ext}}$ and $\delta_{\text{int}}$ are extrinsic and intrinsic reward ratios, $\gamma_{\text{exp}}$ is the discount factor, $\lambda^{\text{ent}}_{\text{exp}}$ is the exploration entropy coefficient adjusted automatically by SAC, and $\lambda^{\text{kld}}_{\text{exp}}$ is the KLD coefficient. Also, note that Eq. (A.3) provides a parameterized and detailed reformulation of Eq. (3) from Section 4.2, explicitly incorporating parameterization and loss scaling details.
Prioritized Skill Refinement with Maximum Return Relabeling
To extract better trajectories and learn skills that effectively solve tasks, the online buffer $\mathcal{B}^{i}_{\text{on}}$ selectively stores high-return trajectories collected during the meta-training phase through the execution of the low-level policy $\pi_{l,\phi}$ and the exploration policy $\pi_{\text{exp},\psi}$. A trajectory $\tau^{i}$ is added to $\mathcal{B}^{i}_{\text{on}}$ if its return $G(\tau^{i})$ exceeds the minimum return in the buffer, $\min_{\tau^{\prime}\in\mathcal{B}^{i}_{\text{on}}}G(\tau^{\prime})$. To refine skills, maximum return relabeling is applied using the parameterized reward model $\hat{R}_{\zeta}$ with parameter $\zeta$. The reward model is trained by minimizing the following MSE loss:
$\mathcal{L}_{\text{reward}}(\zeta):=\mathbb{E}_{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\sim\mathcal{B}^{i}_{\text{exp}}\cup\mathcal{B}^{i}_{\text{on}}}\Bigl{[}\Bigl{(}\hat{R}_{\zeta}(s^{i}_{t},a^{i}_{t},i)-r^{i}_{t}\Bigr{)}^{2}\Bigr{]}.$ (A.4)
For skill refinement, the low-level skill policy $\pi_{l,\phi}$, skill encoder $q_{\phi}$, and skill prior $p_{\phi}$ are optimized using the following loss function. This incorporates both high-return trajectories from the online buffer $\mathcal{B}^{i}_{\text{on}}$ and trajectories from the offline buffer $\mathcal{B}_{\text{off}}$, weighted by their importance:
$\mathcal{L}_{\text{skill}}(\phi):=(1-\beta)\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim P_{\mathcal{B}_{\text{off}}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl{[}\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr{]}+\frac{\beta}{N_{\mathcal{T},\text{train}}}\sum_{i}\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}^{i}_{\text{on}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl{[}\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr{]},$ (A.5)
where $\beta$ is the buffer control factor defined in Eq. (7), and $\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)$ is the skill learning objective defined in Eq. A.1 for optimizing $\pi_{l,\phi}$, $q_{\phi}$, and $p_{\phi}$. During training, we update the low-level policy $\bar{\pi}_{l,\phi}$, skill encoder $\bar{q}_{\phi}$, and skill prior $\bar{p}_{\phi}$ used for the skill-based meta-RL every $K_{\text{iter}}$ iterations. Specifically, the updates are performed as follows: $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
Skill-based Meta-RL
To improve skills through prioritized refinement, the updated and fixed low-level skill policy $\bar{\pi}_{l,\phi}$ and skill prior $\bar{q}_{\phi}$ are utilized to train the high-level policy following the SiMPL framework introduced in Section 3. The objective is to select skill representations $z$ that maximize task returns while ensuring the high-level policy remains close to the skill prior for stable and efficient learning. The high-level policy $\pi_{h,\theta}$ and value function $Q_{h,\theta}$ are parameterized by $\theta$ and trained using the soft actor-critic (SAC) framework, with the RL loss functions defined as:
$\mathcal{L}_{h}^{\text{critic}}(\theta):=\mathbb{E}_{\begin{subarray}{c}(s_{t},z_{t},r^{h}_{t},s_{t+H_{s}-1})\sim\mathcal{B}^{\mathcal{T}}_{h},e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})\\ z_{t+1}\sim\pi_{h,\theta}(\cdot|s_{t+H_{s}-1},e^{\mathcal{T}})\end{subarray}}\biggl{[}\frac{1}{2}\biggl{(}&Q_{h,\theta}(s_{t},z_{t},e^{\mathcal{T}})-\Bigl{(}r^{h}_{t}+\gamma_{h}\bigl{(}Q_{h,\theta}(s_{t+H_{s}-1},z_{t+1},e^{\mathcal{T}})\\ &-\lambda^{\text{kld}}_{h}\mathcal{D}_{\text{KL}}\bigl{(}\pi_{h,\theta}(\cdot|s_{t+H_{s}-1},e^{\mathcal{T}})\,\big{|}\big{|}\,\bar{p}_{\phi}(\cdot|s_{t+H_{s}-1})\bigr{)}\bigr{)}\Bigr{)}\biggr{)}^{2}\biggr{]}\end{aligned}\\ &\mathcal{L}_{h}^{\text{actor}}(\theta):=\mathbb{E}_{\begin{subarray}{c}s_{t}\sim\mathcal{B}^{\mathcal{T}}_{h},e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})\\ z_{t}\sim\pi_{h,\theta}(\cdot|s_{t},e^{\mathcal{T}})\end{subarray}}\biggl{[}\lambda^{\text{kld}}_{h}\mathcal{D}_{\text{KL}}\Bigl{(}\pi_{h,\theta}(\cdot|s_{t},e^{\mathcal{T}})\,\Big{|}\Big{|}\,\bar{p}_{\phi}(\cdot|s_{t})\Bigr{)}-Q_{h,\theta}(s_{t},z_{t},e^{\mathcal{T}})\biggr{]},\end{aligned}$ (A.6)
where $q_{e,\theta}$ is the parameterized task encoder with parameter $\theta$, $\gamma_{h}$ is the high-level discount factor, and $\lambda^{\text{kld}}_{h}$ is the high-level KLD coefficient. The term $r^{h}_{t}=\sum^{t+H_{s}-1}_{k=t}r_{k}$ represents the cumulative rewards, with states and rewards obtained by executing the low-level policy $\bar{\pi}_{l,\phi}$ using $z_{t}\sim\pi_{h,\theta}(\cdot|s_{t})$ over $H_{s}$ timesteps. The context $c^{\mathcal{T}}={(s_{k},z_{k},r_{k}^{h},s_{k+H_{s}-1})}^{N_{\text{prior}}}_{k=1}$, where $N_{\text{prior}}$ is the number of context transitions, denotes the high-level transition set of task $\mathcal{T}$. This context is used to select the task representation $e^{\mathcal{T}}$ from the task encoder $q_{e,\theta}$. Also, note that Eq. (A.6) is a parameterized modification of Eq. (2) from Section 3.
A.3 Meta-Test Phase
After completing the meta-train phase of PRISM, the meta-test phase is performed on the test task set $\mathcal{M}_{\mathrm{test}}$. In this phase, previously learned components, including the task encoder $q_{e,\theta}$, low-level skill policy $\bar{\pi}_{l,\phi}$, and skill prior $\bar{p}_{\phi}$, are kept fixed and are no longer updated. Only the high-level policy $\pi_{h,\theta}$ and high-level value function $Q_{h,\theta}$ are trained for each test task using the soft actor-critic (SAC) framework.
During meta-testing, for each test task $\mathcal{T}$, the task representation $e^{\mathcal{T}}$ is inferred from the fixed task encoder $q_{e,\theta}$. The SAC algorithm is then applied to optimize the high-level policy and value function for the specific test task, following the same loss functions as defined in Eq. (A.6) from the meta-training phase. This approach ensures efficient adaptation to unseen tasks by leveraging the fixed, pre-trained low-level skills and task representations.
Algorithm 1 PRISM: Meta-Train Phase
0:  Training tasks $\mathcal{M}_{\text{train}}$, offline dataset $\mathcal{B}_{\text{off}}$, low-level policy $\pi_{l,\phi}$, skill encoder $q_{\phi}$, and skill prior $p_{\phi}$.
0:  High-level policy $\pi_{h,\theta}$, exploration policy $\pi_{\text{exp},\psi}$, task encoder $q_{e,\theta}$, reward model $\hat{R}_{\zeta}$, and value functions $Q_{h,\theta}$, $Q_{\text{exp},\psi}$.
1:  Fix $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
2:  for iteration $k=1,2,~{}\cdots$ do
3:     for task $i=1$ to $N_{\mathcal{T},\text{train}}$ do
4:        Collect high-level trajectories $\tau_{h}^{i}$ and low-level trajectories $\tau_{l}^{i}$ using $\pi_{h,\theta}$ with $\bar{\pi}_{l,\phi}$, $q_{e,\theta}$.
5:        Collect exploration trajectories $\tau_{\text{exp}}^{i}$ using $\pi_{\text{exp},\psi}$.

### v3

@@ 7 Conclusion
In this paper, we propose SISL, a robust skill-based meta-RL framework designed to address noisy offline demonstrations in long-horizon tasks. Through self-improving skill refinement and prioritization via maximum return relabeling, SISL effectively prioritizes task-relevant trajectories for skill learning and enabling efficient exploration and targeted skill optimization. Experimental results highlight its robustness to noise and superior performance across various environments, demonstrating its potential for scalable meta-RL in real-world applications where data quality is critical.
Ethics statement
By introducing self-improving skill refinement and skill prioritization via maximum return relabeling, SISL improves the stability and generalizability of skill learning, allowing agents to adapt rapidly to new tasks even when data quality is imperfect. This advancement has the potential to make reinforcement learning more practical and reliable in real-world settings where collecting high-quality data is difficult or expensive. While this framework may have potential societal implications by enhancing reinforcement learning’s real-world applicability, we believe it is primarily foundational in nature and does not introduce any new risks of malicious use.
Reproducibility statement
To reproduce SISL, we provide the loss function redefined as neural network parameters, along with the meta-train and meta-test algorithm tables in Appendix B. For the algorithm implementation, we provide the system specifications used for experiments, the source of the baseline algorithm, environment details, the offline dataset construction method, and the hyperparameter setup in Appendix C. Additionally, we provide the anonymized code for SISL in the supplementary material, enabling the reproduction of the proposed algorithm and experiment results.
Appendix A The use of Large Language Models (LLMs)
We wrote the entire manuscript ourselves, including the main text and the appendix. We used large language models only for copy editing to improve spelling and readability, and we verified all suggested revisions before incorporation. LLMs were not used to generate ideas, methods, analyses, results, code, figures, or citations beyond minor edits. All technical content and experiments were conceived, implemented, and validated by the authors. We manually audited every citation and numerical claim and accept full responsibility for the manuscript.
Appendix B Implementation Details on SISL
This section provides a detailed implementation of the proposed SISL framework. As outlined in Section 4, SISL begins with an initial skill learning phase (pre-train) to train the low-level skill policy, skill encoder, and skill prior. It then progresses to the meta-train phase, where decoupled policy learning is performed using high-level policy, task encoder, and skill-improvement policy. Also, self-improvement skill learning is executed via maximum return relabeling using reward model, low-level skill policy, skill encoder, and skill prior. Finally, in the meta-test phase, rapid adaptation to the target task is achieved via fine-tuning based on the trained high-level policy and task encoder. Section B.1 details the initial skill learning phase, Section B.2 elaborates on the meta-train phase, and Section B.3 explains the meta-test phase. All loss functions in SISL are redefined in terms of the neural network parameters of its policies and models. Additionally, the overall structure for the meta-train and meta-test phases is provided in Algorithms 1 and 2.
B.1 Initial Skill Learning Phase
@@
$\begin{aligned} &\mathcal{L}_{\text{spirl}}(\phi)\\ &:=\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}_{\text{off}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl[\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr]\\ &\ \ \begin{aligned} =\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}_{\text{off}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Biggl[-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l,\phi}(a_{k}|s_{k},z)&+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\Bigl(q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\,\Big|\Big|\,\mathcal{N}(\mathbf{0},\mathbf{I})\Bigr)\\ &+\mathcal{D}_{\text{KL}}\Bigl(\lfloor q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\rfloor\,\Big|\Big|\,p_{\phi}(\cdot|s_{t})\Bigr)\Biggr],\end{aligned}\end{aligned}$
where $\lfloor\cdot\rfloor$ represents the stop gradient operator, which prevents the KL term for skill prior learning from influencing the skill encoder. Using the pre-trained $\pi_{l,\phi}$, $q_{\phi}$, and $p_{\phi}$, SISL refines skills during the meta-train phase to further enhance task-solving capabilities.
B.2 Meta-Train Phase
As described in Section 4, SISL comprises two main processes: Decoupled Policy Learning, which explores task-relevant behavior near skill distribution using skill-improvement policy, and trains the high-level policy, task encoder to effectively utilize learned skills for solving tasks; and Self-Improvement Skill Learning, which improves skills using prioritization via maximum return relabeling. Detailed explanations for each process are provided below.
Decoupled Policy Learning
As described in Section 4.2, the skill-improvement policy $\pi_{\text{imp},\psi}$, parameterized by $\psi$, is designed to expand the skill distribution and discover task-relevant behaviors near trajectories stored in the prioritized on-policy buffer $\mathcal{B}_{\text{on}}^{i}$ for each training task $\mathcal{T}^{i}$. This buffer prioritizes trajectories that best solve the tasks. Additionally, the state-action value function $Q_{\text{imp},\psi}$, also parameterized by $\psi$, is defined to train the skill-improvement policy using soft actor-critic (SAC). To enhance exploration, both extrinsic reward $r_{t}^{i}$ and intrinsic reward $r_{\text{int},t}^{i}$ are employed. The intrinsic reward, based on random network distillation (RND), is computed as the L2 loss between a randomly initialized target network $\hat{f}_{\bar{\eta}}^{i}$ and a prediction network $f_{\eta}^{i}$, parameterized by $\eta$ and $\bar{\eta}$, respectively, and is expressed as:
$r_{\text{int},t}^{i}:=\left\|f_{\eta}^{i}(s_{t+1})-\hat{f}_{\bar{\eta}}^{i}(s_{t+1})\right\|^{2}_{2},$ (B.2)
where $i$ is the task index, and $f_{\eta}$ is updated to minimize this loss. A dropout layer is applied to $f_{\eta}$ to prevent over-sensitivity to state $s$. The RL loss functions of SAC for training the skill-improvement policy $\pi_{\text{imp},\psi}$ and the state-action value function $Q_{\text{imp},\psi}$ using the intrinsic reward $r_{\text{int},t}$ are defined as follows:
$\begin{aligned} &\begin{aligned} \mathcal{L}_{\text{imp}}^{\text{critic}}(\psi):=\sum_{i}\mathbb{E}_{\begin{subarray}{c}(s_{t},a_{t},r_{t}^{i},s_{t+1})\sim\mathcal{B}^{i}_{\text{imp}}\cup\mathcal{B}^{i}_{\text{on}}\\ a_{t+1}\sim\pi_{\text{imp},\psi}(\cdot|s_{t},i)\end{subarray}}\biggl[\frac{1}{2}\biggl(Q_{\text{imp},\psi}(s_{t},a_{t},i)-\Bigl(\delta_{\text{ext}}r_{t}^{i}+&\delta_{\text{int}}r_{\text{int,t}}^{i}+\gamma_{\text{imp}}\bigl(Q_{\text{imp},\psi}(s_{t+1},a_{t+1},i)\\ &+\lambda^{\text{ent}}_{\text{imp}}\log\pi_{\text{imp},\psi}(a_{t+1}|s_{t+1},i)\bigr)\Bigr)\biggr)^{2}\biggr]\end{aligned}\\ &\begin{aligned} \mathcal{L}_{\text{imp}}^{\text{actor}}(\psi):=&\sum_{i}\mathbb{E}_{\begin{subarray}{c}s_{t}\sim\mathcal{B}^{i}_{\text{imp}}\cup\mathcal{B}^{i}_{\text{on}}\\ a_{t}\sim\pi_{\text{imp},\psi}(\cdot|s_{t},i)\end{subarray}}\biggl[\lambda^{\text{ent}}_{\text{imp}}\log\pi_{\text{imp},\psi}(a_{t}|s_{t},i)-Q_{\text{imp},\psi}(s_{t},a_{t},i)\biggr]-\lambda^{\text{kld}}_{\text{imp}}\sum_{i}\mathbb{E}_{(s_{t},a_{t})\sim\mathcal{B}^{i}_{\text{on}}}\biggl[\log\pi_{\text{imp},\psi}(a_{t}|s_{t},i)\biggr].\end{aligned}\end{aligned}$
Here, $\delta_{\text{ext}}$ and $\delta_{\text{int}}$ are extrinsic and intrinsic reward ratios, $\gamma_{\text{imp}}$ is the discount factor, $\lambda^{\text{ent}}_{\text{imp}}$ is the exploration entropy coefficient adjusted automatically by SAC, and $\lambda^{\text{kld}}_{\text{imp}}$ is the KLD coefficient. Also, note that Eq. (B.3) provides a parameterized and detailed reformulation of Eq. (3) from Section 4.2, explicitly incorporating parameterization and loss scaling details.
To mutually update skill selection based on the refined skills, the updated and fixed low-level skill policy $\bar{\pi}_{l,\phi}$ and skill prior $\bar{q}_{\phi}$ are utilized to train the high-level policy following the SiMPL framework introduced in Section 3. The objective is to select skill representations $z$ that maximize task returns while ensuring the high-level policy remains close to the skill prior for stable and efficient learning. The high-level policy $\pi_{h,\theta}$ and value function $Q_{h,\theta}$ are parameterized by $\theta$ and trained using the soft actor-critic (SAC) framework, with the RL loss functions defined as:
$\begin{aligned} &\begin{aligned} \mathcal{L}_{h}^{\text{critic}}(\theta):=\mathbb{E}_{\begin{subarray}{c}(s_{t},z_{t},r^{h}_{t},s_{t+H_{s}-1})\sim\mathcal{B}^{\mathcal{T}}_{h},e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})\\ z_{t+1}\sim\pi_{h,\theta}(\cdot|s_{t+H_{s}-1},e^{\mathcal{T}})\end{subarray}}\biggl[\frac{1}{2}\biggl(&Q_{h,\theta}(s_{t},z_{t},e^{\mathcal{T}})-\Bigl(r^{h}_{t}+\gamma_{h}\bigl(Q_{h,\theta}(s_{t+H_{s}-1},z_{t+1},e^{\mathcal{T}})\\ &-\lambda^{\text{kld}}_{h}\mathcal{D}_{\text{KL}}\bigl(\pi_{h,\theta}(\cdot|s_{t+H_{s}-1},e^{\mathcal{T}})\,\big|\big|\,\bar{p}_{\phi}(\cdot|s_{t+H_{s}-1})\bigr)\bigr)\Bigr)\biggr)^{2}\biggr]\end{aligned}\\ &\mathcal{L}_{h}^{\text{actor}}(\theta):=\mathbb{E}_{\begin{subarray}{c}s_{t}\sim\mathcal{B}^{\mathcal{T}}_{h},e^{\mathcal{T}}\sim q_{e,\theta}(\cdot|c^{\mathcal{T}})\\ z_{t}\sim\pi_{h,\theta}(\cdot|s_{t},e^{\mathcal{T}})\end{subarray}}\biggl[\lambda^{\text{kld}}_{h}\mathcal{D}_{\text{KL}}\Bigl(\pi_{h,\theta}(\cdot|s_{t},e^{\mathcal{T}})\,\Big|\Big|\,\bar{p}_{\phi}(\cdot|s_{t})\Bigr)-Q_{h,\theta}(s_{t},z_{t},e^{\mathcal{T}})\biggr],\end{aligned}$
where $q_{e,\theta}$ is the parameterized task encoder with parameter $\theta$, $\gamma_{h}$ is the high-level discount factor, and $\lambda^{\text{kld}}_{h}$ is the high-level KLD coefficient. The term $r^{h}_{t}=\sum^{t+H_{s}-1}_{k=t}r_{k}$ represents the cumulative rewards, with states and rewards obtained by executing the low-level skill policy $\bar{\pi}_{l,\phi}$ using $z_{t}\sim\pi_{h,\theta}(\cdot|s_{t})$ over $H_{s}$ timesteps. The context $c^{\mathcal{T}}={(s_{k},z_{k},r_{k}^{h},s_{k+H_{s}-1})}^{N_{\text{prior}}}_{k=1}$, where $N_{\text{prior}}$ is the number of context transitions, denotes the high-level transition set of task $\mathcal{T}$. This context is used to select the task representation $e^{\mathcal{T}}$ from the task encoder $q_{e,\theta}$. Also, note that Eq. (B.4) is a parameterized modification of Eq. (2) from Section 3.
Self-Improvement Skill Learning
To extract better trajectories and learn skills that effectively solve tasks, the online buffer $\mathcal{B}^{i}_{\text{on}}$ selectively stores high-return trajectories collected during the meta-training phase through the execution of the low-level policy $\pi_{l,\phi}$ and the skill-improvement policy $\pi_{\text{imp},\psi}$. A trajectory $\tau^{i}$ is added to $\mathcal{B}^{i}_{\text{on}}$ if its return $G(\tau^{i})$ exceeds the minimum return in the buffer, $\min_{\tau^{\prime}\in\mathcal{B}^{i}_{\text{on}}}G(\tau^{\prime})$. To refine skills, maximum return relabeling is applied using the parameterized reward model $\hat{R}_{\zeta}$ with parameter $\zeta$. The reward model is trained by minimizing the following MSE loss:
$\mathcal{L}_{\text{reward}}(\zeta):=\mathbb{E}_{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\sim\mathcal{B}^{i}_{\text{imp}}\cup\mathcal{B}^{i}_{\text{on}}}\Bigl[\Bigl(\hat{R}_{\zeta}(s^{i}_{t},a^{i}_{t},i)-r^{i}_{t}\Bigr)^{2}\Bigr].$ (B.5)
For skill learning, the low-level skill policy $\pi_{l,\phi}$, skill encoder $q_{\phi}$, and skill prior $p_{\phi}$ are optimized using the following loss function. This incorporates both high-return trajectories from the online buffer $\mathcal{B}^{i}_{\text{on}}$ and trajectories from the offline buffer $\mathcal{B}_{\text{off}}$, weighted by their importance:
$\mathcal{L}_{\text{skill}}(\phi)$= $\ (1-\beta)\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim P_{\mathcal{B}_{\text{off}}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl[\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr]$ (B.6)
$+\frac{\beta}{N_{\mathcal{T},\text{train}}}\sum_{i}\mathbb{E}_{\begin{subarray}{c}(s_{t:t+H_{s}},a_{t:t+H_{s}})\sim\mathcal{B}^{i}_{\text{on}}\\ z\sim q_{\phi}(\cdot|s_{t:t+H_{s}},a_{t:t+H_{s}})\end{subarray}}\Bigl[\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)\Bigr],$
where $\beta$ is the mixing coefficient defined in Eq. (7), and $\mathcal{L}(\pi_{l,\phi},q_{\phi},p_{\phi},z)$ is the skill learning objective defined in Eq. (B.1) for optimizing $\pi_{l,\phi}$, $q_{\phi}$, and $p_{\phi}$. During training, we update the low-level policy $\bar{\pi}_{l,\phi}$, skill encoder $\bar{q}_{\phi}$, and skill prior $\bar{p}_{\phi}$ used for the skill-based meta-RL every $K_{\text{iter}}$ iterations. Specifically, the updates are performed as follows: $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
B.3 Meta-Test Phase
After completing the meta-train phase of SISL, the meta-test phase is performed on the test task set $\mathcal{M}_{\mathrm{test}}$. In this phase, previously learned components, including the task encoder $q_{e,\theta}$, low-level skill policy $\bar{\pi}_{l,\phi}$, and skill prior $\bar{p}_{\phi}$, are kept fixed and are no longer updated. Only the high-level policy $\pi_{h,\theta}$ and high-level value function $Q_{h,\theta}$ are trained for each test task using the soft actor-critic (SAC) framework.
During meta-testing, for each test task $\mathcal{T}$, the task representation $e^{\mathcal{T}}$ is inferred from the fixed task encoder $q_{e,\theta}$. The SAC algorithm is then applied to optimize the high-level policy and value function for the specific test task, following the same loss functions as defined in Eq. (B.4) from the meta-training phase. This approach ensures efficient adaptation to unseen tasks by leveraging the fixed, pre-trained low-level skills and task representations.
1
Require: Training tasks $\mathcal{M}_{\text{train}}$, offline dataset $\mathcal{B}_{\text{off}}$, low-level policy $\pi_{l,\phi}$, skill encoder $q_{\phi}$, and skill prior $p_{\phi}$.
Initialize: High-level policy $\pi_{h,\theta}$, skill-improvement policy $\pi_{\text{imp},\psi}$, task encoder $q_{e,\theta}$, reward model $\hat{R}_{\zeta}$, and value functions $Q_{h,\theta}$, $Q_{\text{imp},\psi}$.
2 Fix $\bar{\pi}_{l,\phi}\leftarrow\pi_{l,\phi}$, $\bar{q}_{\phi}\leftarrow q_{\phi}$, and $\bar{p}_{\phi}\leftarrow p_{\phi}$.
@@ B.3 Meta-Test Phase
4 for task $i=1$ to $N_{\mathcal{T},\text{train}}$ do
5      Collect high-level trajectories $\tau_{h}^{i}$ and low-level trajectories $\tau_{l}^{i}$ using $\pi_{h,\theta}$ with $\bar{\pi}_{l,\phi}$, $q_{e,\theta}$.
6      Collect skill-improvement trajectories $\tau_{\text{imp}}^{i}$ using $\pi_{\text{imp},\psi}$.